# Optical Entropy and Generalized Thermodynamics of Solitonic Event Horizons

### v3

@@ Abstract
The realization of Hawking radiation in optical analogs has historically focused on kinematic observables, such as the effective temperature determined by the horizon’s surface gravity. A complete thermodynamic description, however, necessitates a rigorous definition of entropy and irreversibility, which has remained elusive in Hamiltonian optical systems. In this work, we bridge this gap by introducing an operational entropy for solitonic event horizons, derived from the spectral partitioning of the optical field into coherent solitonic and incoherent radiative subsystems. We demonstrate that the emission of resonant radiation—mediated by the breaking of soliton integrability due to higher-order dispersion—serves as a fundamental mechanism for entropy production. Numerical simulations of the generalized nonlinear Schrödinger equation confirm that this process satisfies a generalized second law, $\Delta S_{\mathrm{tot}}\geq 0$. These results establish optical event horizons as consistent nonequilibrium thermodynamic systems, offering a new pathway to explore the information-theoretic aspects of analog gravity in laboratory settings.
@@ 3.1 Simulation Parameters
We simulate a physically realizable regime characteristic of supercontinuum generation in photonic crystal fibers [9]. While the equations are solved in dimensionless form to isolate universal behavior, we present results in physical units for experimental clarity. The fiber parameters are chosen to match standard anomalous dispersion profiles: $\beta_{2}=-15\,\mathrm{ps^{2}/km}$, $\beta_{3}=0.1\,\mathrm{ps^{3}/km}$, and $\gamma=0.1\,\mathrm{W^{-1}km^{-1}}$. The input pulse duration is $T_{0}=50\,\mathrm{fs}$ with a soliton order of $N=3.5$, ensuring the system evolves deep into the fission regime where resonant radiation is efficiently generated [14].

### v4

# Generalized Thermodynamics of Solitonic Event Horizons in Dispersive Field Theories

@@ Abstract
The realization of Hawking radiation in optical analogs has historically focused on kinematic observables, such as the effective temperature determined by the horizon’s surface gravity. A complete thermodynamic description, however, necessitates a rigorous definition of entropy and irreversibility, which has remained elusive in Hamiltonian optical systems. In this work, we bridge this gap by introducing an operational entropy for solitonic event horizons, derived from the spectral partitioning of the optical field into coherent solitonic and incoherent radiative subsystems. We demonstrate that the emission of resonant radiation—mediated by the breaking of soliton integrability due to higher-order dispersion—serves as a fundamental mechanism for entropy production. Numerical simulations of the generalized nonlinear Schrödinger equation confirm that this process satisfies a generalized second law, $\Delta S_{\mathrm{tot}}\geq 0$. These results establish event horizons in dispersive field theories as consistent nonequilibrium thermodynamic systems, offering a new pathway to explore the information-theoretic aspects of analog gravity in laboratory settings.
@@ 1 Introduction
From the standpoint of gravitational physics, the defining conceptual role of an event horizon is not associated with a particular microscopic realization, but with the emergence of causal separation and irreversible information flow across a null boundary. This viewpoint is formalized in black hole thermodynamics through the generalized second law (GSL), which asserts that the sum of horizon entropy and exterior entropy is non-decreasing under physically admissible processes [5, 6, 31]. While the geometric area law linking entropy to horizon area is specific to Einstein gravity [4], the logical content of the GSL is more general, relying on coarse-graining over inaccessible degrees of freedom and on the existence of a horizon that mediates information exchange between subsystems. This structural interpretation has been emphasized in studies of horizon entropy as entanglement entropy [8, 25] and underlies the expectation that generalized second-law behavior is not intrinsically tied to the microscopic details of quantum gravity.
Within this broader perspective, the solitonic event horizon investigated here constitutes a representative example of a dispersive, nonequilibrium horizon arising in an integrability-broken field theory. Such horizons are characterized by a localized background that induces mode conversion, a continuum of outgoing dispersive channels that effectively function as an environment, and a dynamical mechanism for irreversible spectral redistribution [29, 30, 21]. In optical systems, these ingredients naturally emerge in the interaction of solitons with higher-order dispersion, leading to resonant radiation and broadband mode coupling [2, 11, 24, 16]. The entropy introduced in this work should therefore be interpreted as an operational, coarse-grained quantity rather than a microscopic state-counting entropy, analogous to entanglement- or environment-induced entropy in open systems [9]. The observed monotonic increase of the total entropy does not depend on quantum statistics or gravitational dynamics, but on the universal interplay between horizons, coarse-graining, and irreversibility. As a result, the conclusions are expected to apply generically to analog horizon systems, including Bose–Einstein condensates and hydrodynamic flows [14, 32, 26], and to inform the interpretation of generalized second-law behavior beyond the specific optical realization considered.
@@ 2 Thermodynamic Formalism
Operationally, this construction is equivalent to tracing over dispersive modes outside a resolution-limited observational algebra, yielding an effective reduced description analogous to entanglement entropy in quantum field theory on curved spacetime [8, 31].
@@ 3.1 Simulation Parameters
We simulate a physically realizable regime characteristic of supercontinuum generation in solid-core photonic crystal fibers. Such systems allow for precise dispersion engineering, where the zero-dispersion wavelength (ZDW) and nonlinear coefficient can be tailored via core geometry to ensure robust soliton fission [11, 33]. While the equations are solved in dimensionless form to isolate universal behavior, we present results in physical units for experimental clarity. The fiber parameters are chosen to match standard anomalous dispersion profiles: $\beta_{2}=-15\,\mathrm{ps^{2}/km}$, $\beta_{3}=0.1\,\mathrm{ps^{3}/km}$, and $\gamma=0.1\,\mathrm{W^{-1}km^{-1}}$. The input pulse duration is $T_{0}=50\,\mathrm{fs}$ with a soliton order of $N=3.5$, ensuring the system evolves deep into the fission regime where resonant radiation is efficiently generated [16].